OpenStudy (anonymous):

use standard form the equation The line with m = -3/2 and b = 6

OpenStudy (anonymous):

Standard form is : y=mx+d Just get -3/2 to m and 6 to b . Just this .:)

OpenStudy (anonymous):

how should i find the x and y????

OpenStudy (uri):

m is slope... So, y=mx+b y=3/2x +6

OpenStudy (anonymous):

They are in the form ! We write them :)

OpenStudy (anonymous):

We dont give numbers for them . y=5x+1 Means the line s points can be : 26=5x+1 x=5 (5,26) Got it ?:)

OpenStudy (anonymous):

so basically the "m" is the "x" and the "b" is the "Y" in standard form

OpenStudy (uri):

basically 3/2 x is y intercept.

OpenStudy (anonymous):

ok get it thankx guys..:)

OpenStudy (anonymous):

Welcome .